# Balancing Act: Prioritization Strategies for LLM-Designed Restless Bandit Rewards

### v2

@@ C.1 Whittle Index Policy
To implement the Whittle Index computation, we use the method in (Qian et al. 2016b) based on binary search. Additionally, for all the experiments, we cache computed whittle index for a given set of transition probabilities and reward functions to reduce computation time.
C.2 DLM Pipeline
In our work, we use a modified version of the DLM pipeline (Behari et al. 2024), which employs the Whittle Index policy as the planner for RMAB. Our approach differs from Behari et al. (2024) in that we use the Whittle Index policy specifically for simulating RMAB, whereas Behari et al. (2024) use the PreFeRMAB policy (Zhao and Behari et al. 2023). This modification allows for faster and more stable simulations and effectively decouples the learning problem from the planning problem.
@@ C.2 DLM Pipeline
C.3 Synthetic Dataset Generation
@@
Table 2: Weight vector parameters for different synthetic datasets.
@@
C.4 Real World Dataset
As described in Section 6.1, one challenge in estimating transition probabilities under active actions is the rarity of such transitions due to the limited number of service calls. Consequently, the estimates for active transition probabilities tend to be noisy. To mitigate this issue, we follow a procedure similar to that in the synthetic domain to construct the datasets, but with some modifications. Specifically, instead of uniformly sampling passive probabilities in step 1, we use the transition probabilities estimated from real-world data (Verma et al. 2023). Next, in step 2, we use the attributes describing the real-world beneficiaries to define features. Thus, instead of features A, B and C, we use features Age, Income and Education. Finally, we use the same weight vector values as in Table 2 to generate three real world datasets. This allows us to generate multiple datasets with varying levels of effect of intervention based on features while still using the realistic passive transition probabilities and feature distributions.
C.5 Hyperparameters
@@ C.6 Computing Resources
Appendix D Additional Results
D.1 Composite Prioritization Prompts
@@
Table 3: Results summary comparing different reward function choice strategies aggregated across the three synthetic Datasets. Cells in bold indicate the top 2 best values (higher is better). DLM: Decision-Language Model, SCLM: Social Choice Language Model, SCLM-SIM: Simulation based Scorer Model, SCLM-LLM: LLM based Scorer Model
@@
Table 4: Results summary comparing different reward function choice strategies aggregated across the three Real World Datasets. Cells in bold indicate the top 2 best values (higher is better). DLM: Decision-Language Model, SCLM: Social Choice Language Model, SCLM-SIM: Simulation based Scorer Model, SCLM-LLM: LLM based Scorer Model
@@
Table 5: Results comparing different reward function choice selection strategies for deciding tradeoff between utility maximization and preference alignment, aggregated across the three synthetic datasets (left) and three real world datasets (right). Social Choice Language Model has the highest percent change in reward for the desired attribute while also ensuring the low drop in utility as compared to default.
Synthetic Real World
% Change in Desired Attribute Utility Metrics % Change in Desired Attribute Utility Metrics
Method top/bottom 1 bucket top/bottom 2 buckets Normalized Utility Score % drop in Utility top/bottom 1 bucket top/bottom 2 buckets Normalized Utility Score % drop in Utility
DLM-PrioritizationOnly -0.812$\pm$0.85 0.072$\pm$0.83 0.507$\pm$0.01 -4.867$\pm$0.21 5.678$\pm$0.67 1.177$\pm$0.43 0.327$\pm$0.01 -6.051$\pm$0.14
DLM-ExtendedPrompt -2.508$\pm$1.2 -1.089$\pm$1.26 0.548$\pm$0.02 -3.825$\pm$0.29 3.416$\pm$1.04 0.461$\pm$0.64 0.347$\pm$0.02 -5.689$\pm$0.25
SCLM 12.389$\pm$1.01 10.276$\pm$0.8 0.71$\pm$0.01 -1.141$\pm$0.16 9.058$\pm$0.61 4.868$\pm$0.4 0.532$\pm$0.01 -3.662$\pm$0.16
D.2 Effect of Social Choice Function
Figure 5: Comparison of different methods using the prompt “Prioritize agents with a low value for feature $B$ and agents with a low value for feature $C$”.
While both the Utilitarian and Egalitarian social welfare function leads to good overall results, the choice between them significantly influences which reward function is selected at the instance level. Figure 5 shows one example instance for a prompt with two prioritization clauses. As in Figure 1, each point corresponds to a candidate reward function with the axes measuring alignment with the two respective clauses. (as described in the beginning of the section 6.3). The SCLM-Utilitarian model chooses a reward function from the Pareto frontier that shows a much stronger effect for the second clause (i.e., the utility increase for agents with a low value of feature $C$ is more pronounced). In contrast, the reward function selected by SCLM-Egalitarian is much more balanced. DLM-based baselines fail to pick a reward function from the Pareto frontier.
D.3 Tradeoffs in preferences and cumulative utility
To show the general applicability of our framework, we consider prompts where the user can specify preferences for multiple attributes and also specify a preference clause that prefers maximizing utility. For all 6 singular and 12 composite prompts, we add additional clauses to maximize utility (or minimize drop in utility). As a baseline, we consider two models: DLM prompted only by the prioritization clause (DLM-Prioritization Only) and DLM prompted by both the prioritization clause(s) and a clause requesting the maximization of overall utility (DLM-ExtendedPrompt).
To assess alignment with prioritization clauses, we calculate the average change in utility for the prioritized subgroup, following the method outlined in Section 6.3. To quantify the score for maximizing utility, we use the utility generated by the default reward function, normalizing this value between 0 and 1 to ensure all score components are on the same scale. Finally, with scores for all prompt components, we can apply a user-defined Social Welfare Function.
Table 5 shows the results. Comparing DLM-Prioritization Only and DLM-ExtendedPrompt, we find that adding the additional objective to the prompt does not result in a better performance on both the synthetic and real-world domains. In contrast, SCLM chooses reward functions resulting in significantly higher utility increases for the prioritized subpopulations and significantly higher utility (measured through Normalized Utility Score) and lower drop in utility as compared to default reward function. The fact that SCLM performs advantageously for both (conflicting) objectives highlights the quality of the pipeline and its capabilities to effectively address multiple objectives (of different types).
Appendix E Prompt Texts
@@ 14. Feature C bucket 5 - Binary
Figure 6: Prompt passed to the LLM to generate reward function based on the problem scenarios in the Synthetic Domain.
@@ 11. Education level 7/7 – Post graduate - Binary
Figure 7: Prompt passed to the LLM to generate reward function based on the problem scenarios in the Real World Domain.
@@
Figure 8: Prompt passed to the LLM to choose a reward function based on the context of problem scenario in the Synthetic Domain, the generated reward functions and the reward distribution corresponding to every reward function.
@@
Figure 9: Prompt passed to the LLM to choose a reward function based on the context of problem scenario in Real World Domain, the generated reward functions and the reward distribution corresponding to every reward function.
@@
Figure 10: Enhanced Prompt passed to the LLM to choose a reward function based on the context of the problem scenario in the Real World Domain, the generated reward functions, the reward distributions corresponding to every reward function and additional examples on what to look at when choosing a reward function aligned with the preference.
@@
Figure 11: Enhanced Prompt passed to the LLM to choose a reward function based on the context of the problem scenario in the Real World Domain, the generated reward functions, the reward distributions corresponding to every reward function and additional examples on what to look at when choosing a reward function aligned with the preference.
@@
My goal was to create a Python reward function for RL in resource allocation, with the objective of:
Focus on the young mothers by age and also focus on those with low education. I tried several reward functions for this task. Below, I have the given reward function, and the corresponding distribution of reward achieved across 20 agent features.
@@
Figure 12: Enhanced Prompt passed to the LLM to choose a reward function based on the context of the problem scenario in the Real World Domain, the generated reward functions, the reward distributions corresponding to every reward function and additional information to minimize the unintended utility shifts in dimensions not specified in the preference.
@@
Figure 13: Continued: Enhanced Prompt passed to the LLM to choose a reward function based on the context of the problem scenario in the Real World Domain, the generated reward functions, the reward distributions corresponding to every reward function and additional information to minimize the unintended utility shifts in dimensions not specified in the preference.
@@
My goal was to create a Python reward function for RL in resource allocation, with the objective of:
Focus on the young mothers by age and also focus on those with low education. I tried several reward functions for this task. Below, I have the given reward function, and the corresponding distribution of reward achieved across 20 agent features.
@@
Figure 14: Enhanced Prompt passed to the LLM to choose a reward function based on the context of the problem scenario in the Real World Domain, the generated reward functions, the reward distributions corresponding to every reward function and additional information to maximize the overall utility.
@@
Figure 15: Continued: Enhanced Prompt passed to the LLM to choose a reward function based on the context of the problem scenario in the Real World Domain, the generated reward functions, the reward distributions corresponding to every reward function and additional information to maximize the overall utility.

### v3

@@ B.1. Whittle Index Policy
To implement the Whittle Index computation, we use the method in Qian et al. (2016b) based on binary search. Additionally, for all the experiments, we cache computed whittle index for a given set of transition probabilities and reward functions to reduce computation time.
B.2. DLM Pipeline
In our work, we use a modified version of the DLM pipeline Behari et al. (2024), which employs the Whittle Index policy as the planner for RMAB. Our approach differs from Behari et al. (2024) in that we use the Whittle Index policy specifically for simulating RMAB, whereas Behari et al. (2024) use the PreFeRMAB policy (Zhao and Behari et al. 2023). This modification allows for faster and more stable simulations and effectively decouples the learning problem from the planning problem.
@@ B.2. DLM Pipeline
B.3. Synthetic Dataset Generation
@@
Table 2. Weight vector parameters for different synthetic datasets.
@@
B.4. Real World Dataset
As described in Section 6.1, one challenge in estimating transition probabilities under active actions is the rarity of such transitions due to the limited number of service calls. Consequently, the estimates for active transition probabilities tend to be noisy. To mitigate this issue, we follow a procedure similar to that in the synthetic domain to construct the datasets, but with some modifications. Specifically, instead of uniformly sampling passive probabilities in step 1, we use the transition probabilities estimated from real-world data Verma et al. (2023). Next, in step 2, we use the attributes describing the real-world beneficiaries to define features. Thus, instead of features A, B and C, we use features Age, Income and Education. Finally, we use the same weight vector values as in Table 2 to generate three real world datasets. This allows us to generate multiple datasets with varying levels of effect of intervention based on features while still using the realistic passive transition probabilities and feature distributions.
B.5. Hyperparameters
@@ B.6. Computing Resources
Appendix C Additional Results
C.1. Composite Prioritization Prompts
@@
Table 3. Results summary comparing different reward function choice strategies aggregated across the three synthetic Datasets. Cells in bold indicate the top 2 best values (higher is better). DLM: Decision-Language Model, SCLM: Social Choice Language Model, SCLM-SIM: Simulation based Scorer Model, SCLM-LLM: LLM based Scorer Model
@@
Table 4. Results summary comparing different reward function choice strategies aggregated across the three Real World Datasets. Cells in bold indicate the top 2 best values (higher is better). DLM: Decision-Language Model, SCLM: Social Choice Language Model, SCLM-SIM: Simulation based Scorer Model, SCLM-LLM: LLM based Scorer Model
@@
Table 5. Results comparing different reward function choice selection strategies for deciding tradeoff between preference alignment, unintended shift minimization and utility maximization, aggregated across the three synthetic datasets (left) and three real world datasets (right). A higher summed % change in desired features implies a better alignment, whereas less unintended shift is better.
Real World Synthetic
Method
Summed % Change in
Desired Feature(s)
Unintended
Shift
Summed % Change in
Desired Feature(s)
Unintended
Shift
DLM-PrioritizationOnly 6.809$\pm$0.86 0.302$\pm$0.02 0.06$\pm$1.13 1.15$\pm$0.04
DLM-ExtendedPrompt-Fair -0.254$\pm$0.73 0.276$\pm$0.02 -1.882$\pm$1.31 1.173$\pm$0.06
SCLM-PrioritizationOnly 13.131$\pm$0.86 0.316$\pm$0.02 9.349$\pm$1.45 1.218$\pm$0.05
SCLM-ExtendedPrompt-Fair 15.364$\pm$0.94 0.099$\pm$0.01 10.006$\pm$1.13 0.239$\pm$0.02
Table 6. Results comparing different reward function choice selection strategies for deciding tradeoff between preference alignment, unintended shift minimization and utility maximization, aggregated across the three synthetic datasets (left) and three real world datasets (right). A higher summed % change in desired features implies a better alignment, whereas less unintended shift is better.
Real World Synthetic
Method
Summed % Change in
Desired Feature(s)
Normalized
Utility Score
% Drop in
Utility
Summed % Change in
Desired Feature(s)
Normalized
Utility Score
% Drop in
Utility
DLM-PrioritizationOnly 6.809$\pm$0.86 0.317$\pm$0.01 -6.233$\pm$0.17 0.06$\pm$1.13 0.485$\pm$0.01 -5.403$\pm$0.28
DLM-ExtendedPrompt-Util 3.416$\pm$1.04 0.347$\pm$0.02 -5.689$\pm$0.25 -2.508$\pm$1.2 0.548$\pm$0.02 -3.825$\pm$0.29
SCLM-PrioritizationOnly 13.131$\pm$0.86 0.397$\pm$0.01 -5.372$\pm$0.19 9.349$\pm$1.45 0.614$\pm$0.01 -2.8$\pm$0.24
SCLM-ExtendedPrompt-Util 8.869$\pm$0.73 0.519$\pm$0.02 -3.91$\pm$0.21 13.211$\pm$1.31 0.736$\pm$0.01 -0.568$\pm$0.12
SCLM-Full 11.496$\pm$0.87 0.529$\pm$0.01 -3.85$\pm$0.18 9.166$\pm$1.05 0.705$\pm$0.01 -1.044$\pm$0.11
Appendix D Prompt Texts
@@ 14. Feature C bucket 5 - Binary
Figure 6. Prompt passed to the LLM to generate reward function based on the problem scenarios in the Synthetic Domain.
@@ 11. Education level 7/7 – Post graduate - Binary
Figure 7. Prompt passed to the LLM to generate reward function based on the problem scenarios in the Real World Domain.
@@
Figure 8. Prompt passed to the LLM to choose a reward function based on the context of problem scenario in the Synthetic Domain, the generated reward functions and the reward distribution corresponding to every reward function.
@@
Figure 9. Prompt passed to the LLM to choose a reward function based on the context of problem scenario in Real World Domain, the generated reward functions and the reward distribution corresponding to every reward function.
@@
Figure 10. Enhanced Prompt passed to the LLM to choose a reward function based on the context of the problem scenario in the Real World Domain, the generated reward functions, the reward distributions corresponding to every reward function and additional examples on what to look at when choosing a reward function aligned with the preference.
@@
Figure 11. Enhanced Prompt passed to the LLM to choose a reward function based on the context of the problem scenario in the Real World Domain, the generated reward functions, the reward distributions corresponding to every reward function and additional examples on what to look at when choosing a reward function aligned with the preference.
@@
My goal was to create a Python reward function for RL in resource allocation, with the objective of: Focus on the young mothers by age and also focus on those with low education. I tried several reward functions for this task. Below, I have the given reward function, and the corresponding distribution of reward achieved across 20 agent features.
@@
Figure 12. Enhanced Prompt passed to the LLM to choose a reward function based on the context of the problem scenario in the Real World Domain, the generated reward functions, the reward distributions corresponding to every reward function and additional information to minimize the unintended utility shifts in dimensions not specified in the preference.
@@
Figure 13. Continued: Enhanced Prompt passed to the LLM to choose a reward function based on the context of the problem scenario in the Real World Domain, the generated reward functions, the reward distributions corresponding to every reward function and additional information to minimize the unintended utility shifts in dimensions not specified in the preference.
@@
My goal was to create a Python reward function for RL in resource allocation, with the objective of: Focus on the young mothers by age and also focus on those with low education. I tried several reward functions for this task. Below, I have the given reward function, and the corresponding distribution of reward achieved across 20 agent features.
@@
Figure 14. Enhanced Prompt passed to the LLM to choose a reward function based on the context of the problem scenario in the Real World Domain, the generated reward functions, the reward distributions corresponding to every reward function and additional information to maximize the overall utility.
@@
Figure 15. Continued: Enhanced Prompt passed to the LLM to choose a reward function based on the context of the problem scenario in the Real World Domain, the generated reward functions, the reward distributions corresponding to every reward function and additional information to maximize the overall utility.